Use the given data to find the minimum sample size required to estimate the population proportion.
Margin of error: 0.04; confidence level: 95%; from a prior study, p is estimated by the decimal
equivalent of 60%.

Answer by jim_thompson5910(35256)

p = 0.6
z = 1.96 (use a table or calculator)
E = 0.04

------------------------------------

n = p(1-p)(z/E)^2

n = 0.6(1-0.6)(1.96/0.04)^2

n = 576.24

n = 577 ... Round up to the nearest whole number


The min sample size required is 577